OpenStudy (anonymous):

the terminal side of an angle theta in standard position coincides with the line 2x-y=0 in quadrant 3. find cos theta to the nearest ten-thousandth.

iYuko (iyuko):

in quadrant 3 both x and y are negative

iYuko (iyuko):

since (x, y) lies on 2x - y = 0==> 2x - y = 0 ==> y = 2x

iYuko (iyuko):

tan(A) = -y/-x = y/x = 2x/x = 2

iYuko (iyuko):

A = tan^-1(2)

iYuko (iyuko):

equals 63.4349

iYuko (iyuko):

Got it?

OpenStudy (anonymous):

i understand a lot better now. thank you for your help, i really appreciate it n.n